statistik

2.7 An article in The Economist (July 3, 2010) stated that the number of people in prison is 154 per 100,000 in England and Wales, 96 per 100,000 in France, 87 per 100,000 in Germany, and 753 per 100,000 in the United States Explain how to use the relative risk to compare the U.S. rate to the others.
Answer #1

We can find the relative risk of people being in a prison in the US against that in all other countries. For this we will compute the probability of being in prison for all other countries except US and then the probability of being in prison in the US! Then the relative risk for being in prison in the US is:

\(\frac{\text { Probability of being in prison in US }}{\text { Probability of being in prison in all given countries except US }}\)

Find the probability of a person being in person in all states except for the US:

Total number of people in prisons in all countries except for the \(U S=154+96+87=337\)

Total number of people being studied outside \(U S=3^{*} 100,000=300,000\)

THE required probability \(=337 / 300,000\)

Probability of being in prison in the US=753/100,000

Thus, Relative ris

$$ \frac{\frac{753}{100000}}{=\frac{337}{300000}}=\frac{753 \cdot 3}{337}=6.703 $$

Conclusion: An individual is 6.7 times more likely to be in prison in the US than other countries!!

answered by: Buship